Planning Page for Eng/D Nights During May, June, and July




Nights Available

UT May 18 - 0.5 night D

  • D-time twilight until midnight
    • DONE 1.5 hours allocated to LBTI D-science
    • DONE 1.5 hours alloacted to LBTI Eng work (NCPA checks on bright stars)
  • PETS on-sky midnight-morning twilight

UT May 21 - 0.5 night Eng

  • SOUL-COM (LBTI) - twilight until handover (00:11)
  • Engineering Tasks
    • PEPSI - AGW4 checkout 1-2hrs
    • PEPSI-Pol on-sky checks
      • From Ilya:

        Name V Inst Fiber BlueArm RedArm Start MST End Duration Phase PID
        alpha Boo -0.05 POL 200P CD3 6x00:05:00 CD5 6x00:05:00 01:55:00 02:29:18 00:34:18 polstd
        HR 7950 3.78 POL 200P CD1 2x00:05:00 CD4 2x00:05:00 02:31:18 02:42:44 00:11:26 fluxstd
        HR 7950 3.78 POL 200P CD2 2x00:05:00 CD5 2x00:05:00 02:42:44 02:54:10 00:11:26 fluxstd
        HR 7950 3.78 POL 200P CD3 2x00:05:00 CD6 2x00:05:00 02:54:10 03:05:36 00:11:26 fluxstd
        gamma Equ 4.70 POL 200P CD3 6x00:05:00 CD5 6x00:05:00 03:07:36 03:41:54 00:34:18 polstd

    • Other tasks from 03:52 (assuming some recofig)
    • ´╗┐Night Lost to Weather

UT May 29 - up to 0.5 night

  • LBTI has relinquished night from SOUL-COM
  • MODS Checkout highest priority -1.5 hours
  • MODS field rotator check - 1-1.5 hours
  • Any remaining time goes to INAF for science

UT May 31 - 0.5 night D

  • Night center 00:17
  • 1st half D/E-time
    • MODS checkout rollover
      • MODS checkout after re-install 1-1.5hrs - highest priority
        • Collimate on target
        • MODS mask alignment check
          • Stonel1 12 35 46 -01 01 00 - visible 1st half
          • StoneO 21 37 22 00 31 54 - visible 2nd half (contingency for June 2)
        • Take some test data (specstd should be sufficient in all modes)
        • DETXY dither tests
      • MODS field correction test - 1.5 hours
    • LUCI M13 pseudo-speckle test (1hr) - D. Thompson
  • 2nd half LBTI Science

UT June 3 - 0.5 Night D

  • D/E time 1st half
    • Rollover MODS Checkout (highest priority)
    • LUCI M13 pseudo-speckle test (1hr) D. Thompson - rollover if not done earlier
    • LUCI-1 AO tests
  • Split night with PEPSI (PI Woodward) - 2nd half

UT June 11 - 0.5 Night Eng

  • Split night with PEPSI (PI. Starrfield)
  • 1st half D/E
  • 2nd haff AZ PEPSI
  • Night center 00:19 - handover 00:13 if LUCI/MODS first, handover 00:09 if LBC first
  • Moon < 1%
  • Plan:
    • Bin-MODS LS Science - J1201- 2.5hrs (Rothberg)
      • start as soon as possible in twilight
      • thin clouds
      • Seeing < 1.2"
      • Obtained half data requested (lost time due to wildfire particles)
    • SpecPhot Std Dual Grating - 0.2 hrs
    • MODS Field Collection 1-1.5 hrs (Eng)
      • Could not complete - error in IDL Code
    • Possible TMS observations in lieu of MODS (C. Veillet)
    • Backup
      • TMS (LBCs)

UT June 28 - 0.5 night Eng

  • Split night with PETS
  • D-time MODS observations - Rothberg
  • PETS 22:00-05:00 hr local time

UT Jul 12 - 0.5 night D

  • Split night with PETS
  • PETS 21:25-03:00 hr local time


Engineering Tasks Submitted

  • DONE PEPSI AGW4 Checkout 1-2hr - allocated
  • PEPSI Pol on-sky flux/Pol standards 2.5hr - allocated
  • DONE LBTI NCPA tests - 1.5hr - allocated
  • DONE MODS Checkout after re-install - 1-1.5 hrs - allocated
  • DONE MODS AGW-1 Hotspot check
  • DONE MODS Field rotator check
    • needs to be redone - but software errors - POSTPONED UNTIL RESTART
  • DONE MODS 64bit testing on-sky - can be done in parallel with other D or E use of MODS
  • DONE LBC Superfoc sequence - 2 hours - no specific dates/time
  • DONE Pseudo-Speckle Observations AO/ESM - 1.5 hours - no specific dates/times
  • DONE LBC TMS testing (binoc mode) - 3 hours - no specific dates
  • Additional LBC TMS Testing - including a pointing model - 3 hours
  • DONE LBC - trailing test at elevations > 80 deg - 2hr - any conditions
  • Pointing collection - MODS or LUCI - PtStars list - any conditions - backup for all other items
  • Photometric Standards
    • LBC
    • MODS
      • spectrophotometric at differeing airmasses
    • LUCI

Science Programs Submitted

  • DONE Follow-up HOSTS target - LBTI 1.5 hours (approved) - scheduled for UT May 18
  • PDS 456 LUCI-AO/ESM observations - visible ~22hr-03hr - proposal submission pending - for future D night - Kuhn (pending AO release)
  • Nova monitoring - PEPSI -20min (morning twilight)
  • Deep MODS spectroscopy of IMBH candidate - 2.5 hrs (approved) - UT Jun 11 - Rothberg
    • Not completed - 1 hour shutter time + specstd obtained

Poor Weather Programs (Science)

  • Local Mergers (PI Rothberg) LD-2019B-005
    • DONE 1 MODS Target
    • 8 LUCI Targets

Location of Files

  • /home/lbto/Service_Observing/2021A_D_E/May_Jun_Jul_2021 - OBSN

-- BarryRothberg - 19 May 2021
Topic revision: r22 - 25 Jun 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback